For each structure below, determine the number of

For each structure below, determine the number of ? electrons. Indicate whether the structure is aromatic, nonaromatic, or antiaromatic. Assume planarity.
Answer #1
Concepts and reason

The concept used in solving the given problem is based on the Huckel’s rule and estimation of the number π\pi -electrons.

The presence of π\pi -bonds and a lone pair of electrons helps in the determination of the number of electrons and the Huckel’s rule help in the determination of the aromaticity of the compound.

Fundamentals

The basic principles of the Huckel’s rule are as follows:

1. The conjugated and delocalized p-orbital of the compound must consist of (4n+2)π\left( {4n + 2} \right)\pi electrons to be aromatic in nature. Here, n is positive integer.

2. The compound should be planar in nature.

3. The compound must be cyclic with conjugated π\pi electrons that can be from π\pi bonds or lone pairs.

4. If the compound does not follow any one of the above three rules it is non-aromatic.

5. If the compound is cyclic, planar but rather than following 4n+24n + 2 follow 4n rule then it is antiaromatic.

(a.1)

There are 33 conjugated double bonds in the compound therefore, number of π\pi electrons will be 66 .

(a.2)

The compound is cyclic, planar and it follows (4n+2)π\left( {4n + 2} \right)\pi electron rule. The compound is also cyclic and planar therefore, this is aromatic in nature.

(b.1)

There are 3 conjugated double bonds and 1 lone pair in the compound therefore, a number of π\pi electrons will be 8.

(b.2)

The number of π\pi electrons are eight. Therefore, it follows 4n rule. The compound is cyclic and planar therefore, it is anti-aromatic.

[Part b.2]

Part b.2

Ans: Part a.1

The number of π\pi electrons present in the given compound are 66 .